can u please solve these 3 problems or me
-8+5x+x=-74
2x+3y=-15,8x+5y=-25.
pv=nRT

Answer by stanbon(48535)

1. -8+6x=-74
6x = -66
x = -11
2.
1st: 2x+3y = -15
2nd: 8x+5y = -25
To eliminate the "x's" you multiply the 1st by "4"
and then subtract it from the 2nd, as follows:
new 1st: 8x+12y = -60
new 2nd: 0x-7y = 35
Solve for "y":
y = -5
Substitute into the original 1st equation:
2x + 3(-5) = -15
2x + -15 = -15
x = 0
Check x=0 and y=-5 in the original 2nd:
8(0)+5(-5) = -25
-25 = -25
So, the answer is x=0, y= -5
3rd Problem: Don't know what you are asked to do with this.
cheers,
Stan H.